the sneaked store at the mall is having a sale. sneakers that sell $68.00 have a discount of 25%. if you buy them on wednesday you get an additional $5.00. what percent is the total discount if you buy on wednesday?
Answer by rfer(16322)

68*.75=51
51-5=46
46/68=0.676
1.00-0.676=0.324=32.4%